the potential different across the lamp is 4.2v the current in the lamp is then 0.19 a calculate the resistance of the lamp
1
Expert's answer
2018-04-29T13:47:07-0400
Ohm's law - I=V/R→R=V/I , calculate the resistance of the lamp - R=(4.2V)/(0.19A)=22.1 Ohm